# TrendFact: A Benchmark for Explainable Hotspot Perception in Fact-Checking with Natural Language Explanation

### v3

@@ 4 FactISR
4.1 Overview
Inspired by the impressive reasoning capabilities of advanced RLMs like DeepSeek-R1 in complex tasks, we believe that they hold strong potential for addressing the challenges of fact-checking. However, there lack effective methods for conveying fact-checking task content to RLMs in a way that facilitates efficient reasoning. To bridge this gap, we propose FactISR, a framework designed to enhance the general fact-checking abilities of RLMs. An overview of FactISR is shown in Figure 2. It comprises three key components: dynamic evidence augmentation, reasoning adaptation via evidence triangulation, and iterative self-reflection via reward decoding. In the following sections, we introduce each of these components in detail.
4.2 Dynamic Evidence Augmentation
High-quality evidence is fundamental to accurate claim verification. It’s an intuitive way to feed all retrieved evidence directly into the RLM to support the fact-checking. However, lengthy evidences often contain irrelevant or noisy information, which can hinder accurate judgment and lead to redundant reasoning time. To address these issues, we develop the Dynamic Evidence Augmentation (DEA) module. Specifically, DEA iteratively incorporates retrieved evidence into the RLM when specific signals are detected in the model’s output. Figure 10 provides an example of the DEA.
4.3 Reasoning Adaptation via Evidence Triangulation
Given the augmented evidence, it remains challenging for RLMs to fully capture the relationships between the evidence and other key elements, such as the claim, explanation, and label, to make accurate judgments. To this end, we introduce a comprehensive reasoning template based on Evidence Triangulation to enhance the RLM’s understanding and adaptation to fact-checking tasks.
Firstly, we construct the evidence triangulation framework based on the in-depth analysis of human reasoning patterns, derived from hundreds of verification actions performed by our annotation team. This framework comprises: (1) the relevance between the claim and evidence, (2) the consistency of key information extracted from evidence, and (3) the conflict within reasoning process. Then, we build a structured reasoning template based on these factors, as depicted in the six interconnected modules shown in Figure 2. Specifically, the RLM is first prompted to generate an initial verification label based on the relevance assessment. Next, it extracts key information from the evidence and evaluates its consistency with the claim, refining the initial label and producing an explanation. Finally, the RLM reflects on the entire reasoning process to identify any internal conflicts and generate feedback signals for further reasoning.
4.4 Iterative Self Reflection via Reward Decoding
Given the reasoning prompt, the ideal behavior for the RLM is to follow it and engage in iterative reasoning to eliminate conflicts within the reasoning process. However, we observe that the RLM occasionally fails to recognize conflicts, resulting in premature termination. To address this limitation and enhance reasoning continuity, we introduce a reward decoding mechanism that increases the likelihood of the RLM entering the iterative reflection phase, thereby resolving unresolved conflicts (as shown in the right part of Figure 2).
Specifically, we define the probability of the token "yes" as the reward signal and designate a specific token sequence near the end of the reasoning template as the reward trigger. Under this setting, the RLM automatically triggers the reward as reasoning nears completion, receiving a gradually decreasing probability for the next token being "yes". This guides RLM enters into iterative reasoning and allow it to reflect on previous conclusions to generate proper judgment. This approach can be formalized as follows:
$\boldsymbol{x}_{h+k}^{\prime}$ $=\mathrm{RD}(\boldsymbol{x}_{h+k})$
=$=\begin{cases}\boldsymbol{x}_{h+k}+\Delta_{0}\cdot\gamma^{i}\cdot\boldsymbol{R},&\textrm{if }\boldsymbol{x}_{h:h+k-1}=\boldsymbol{S}\\ \boldsymbol{x}_{h+k},&\textrm{otherwise}\end{cases}$ (1)
Where $\boldsymbol{x}_{h+k}^{\prime}$ represents the adjusted token at position $h+k$ in the LLM, while $\boldsymbol{x}_{h+k}$ is the originally generated token. $\boldsymbol{x}_{h:h+k-1}$ denotes the continuous token sequence generated by RLM from position $h$ to $h+k-1$. $\boldsymbol{S}$ is the specific token sequence. When $\boldsymbol{x}_{h:h+k-1}$ matches $\boldsymbol{S}$, we calculate the reward vector $\boldsymbol{R}$ (used to assign rewards to certain affirmative tokens) through the initial reward value $\Delta_{0}$ and the decay factor $\gamma$ (with the value range $10<\gamma<$ 1), with the $i$ means iteration step.
@@ 5.1 Setup
For evidence retrieval task, we choose R@k, where k=1,2,3,5. For verification task, we choose F1-macro, Precision, Recall, and Accuracy. For the explanation generation task, we first employ BLEU-4, ROUGE (including ROUGE-1, ROUGE-2, and ROUGE-L), and BERTScore to assess the quality of generated explanation. Then we introduce two novel metrics: ECS and HCPI. Specifically, ECS evaluates the alignment between the system’s generated explanations and verification results, which jointly considers LLM-based explanation scoring and verification accuracy. HCPI integrates multiple attributes, including hotspot indicators, risk factor, and ECS, to assess a fact-checking system’s ability to detect and respond to high-impact erroneous claims. HCPI encourages models to identify and tag high-impact errors, providing explanations for them. It also penalizes incorrect judgments of non-errors, enhancing applicability to real-world platforms. The detailed formulation information is provided in Appendix F.
@@
In this work, we choose the following types methods as baseline, including RLMs, LLMs and existing fact-checking methods. For RLMs, we select the most advanced QwQ-32B, QwQ-32B-Preview qwe (2024), Qwen3-32B(Think) Yang et al. (2025), and DeepSeek-R1 Guo et al. (2025). For LLMs, we choose GPT-4o OpenAI (2024), DeepSeek-v3 Liu et al. (2024), and Qwen2.5-72B-Instruct qwe (2024), Qwen3-32B(No Think). For fact-checking methods, we select two representative automated methods including PROGRAMFC Pan et al. (2023) and CLAIMDECOMP Chen et al. (2022). For retrieve methods, we choose the following advanced methods including BM-25(without date), BM-25, OpenAI’s text-embedding-ada-002 and bge-m3(dense)Chen et al. (2024).
Experimental Settings.
We conduct experiments on pytorch and $1002\times A$ 100 GPUs. The evaluation for ESC was conducted using GPT-4o, while BERTScore evaluations are conducted on bert-base-chinese. More details are shown in Appendix J.
@@
BM25 -w/o date 35.92 48.11 56.54 65.94
BM25 36.70 49.07 57.28 66.75
text-emb-ada-002 28.35 37.42 42.56 61.93
bge-m3(dense) 49.02 61.48 69.67 78.97
@@
Methods Acc F1 P R
PROGRAM-FC 45.24 44.28 43.34 45.30
CLAIMDECOMP 47.48 46.40 45.32 47.53
QwQ-32B-Preview 56.73 53.30 55.92 57.27
Qwen2.5-72B-instruct 58.64 51.96 58.90 56.28
Qwen3-32B(No think) 61.51 55.01 58.11 58.07
DeepSeek-V3 63.42 57.17 60.35 60.44
GPT-4o 62.29 59.45 60.54 63.05
Qwen3-32B(Think) 70.09 65.20 64.97 67.17
DeepSeek-R1 71.67 64.94 65.31 66.00
QwQ-32B 72.67 66.68 66.96 68.00
FactISR(QwQ-32B) 74.32 67.58 69.08 68.06
FactISR(Qwen3-32B) 72.46 65.01 66.07 65.29
@@
QwQ-32B-Preview 0.4923 0.7689 0.1474 0.7479 0.4544 0.2702 0.3928
Qwen2.5-72B-instruct 0.5321 0.7193 0.2994 0.8163 0.6000 0.4128 0.5446
Qwen3-32B(No think) 0.5172 0.7587 0.2740 0.8166 0.5921 0.3949 0.5389
DeepSeek-V3 0.5718 0.7623 0.2609 0.8058 0.5711 0.3705 0.5182
GPT-4o 0.5655 0.7972 0.2351 0.7934 0.5456 0.3380 0.4794
Qwen3-32B(Think) 0.5679 0.8279 0.2378 0.7962 0.5475 0.3402 0.4897
DeepSeek-R1 0.6032 0.8430 0.2144 0.7833 0.5152 0.3111 0.4519
QwQ-32B 0.6110 0.8355 0.2214 0.7858 0.5237 0.3163 0.4622
FactISR(QwQ-32B) 0.6336 0.8375 0.2185 0.7866 0.5251 0.3198 0.4645
FactISR(Qwen3-32B) 0.6157 0.8268 0.2443 0.8015 0.5604 0.3585 0.5097
@@ 5.2 Main Results
We evaluate the four retrieval methods on their ability to retrieve target gold evidence from the evidence library of TrendFact. As shown in Table 2, the strongest bge-m3 only achieves a moderate level of performance with R@5 less than 80%. This indicates that our unique evidence library construction is capable of collecting challenging evidence for distinguishing original gold evidence, thereby increasing the difficulty of TrendFact.
@@
We perform a comprehensive evaluation of various methods on the verification task of TrendFact, including traditional automatic fact-checking methods, LLMs, and RLMs. The results are presented in Table 3, and can be summarized as follows: First, traditional fact-checking models perform the worst, with all accuracy scores falling below 50%. Second, both LLMs and RLMs achieve better performance, exceeding the 50% threshold. Notably, RLMs outperform LLMs, as they are better equipped to handle the complex reasoning required by many TrendFact samples. However, even the strongest RLM, QwQ-32B, fails to achieve an overall 70% verification performance on TrendFact underscoring the significant challenge posed by the high-quality and reasoning-intensive nature of TrendFact. Furthermore, FactISR contributes to an improvement in the verification performance of RLMs. For example, it improves the performance of the best performing models, QwQ-32B and Qwen3-32B, by 1.65% and 2.37% in accuracy. In addition, the enhanced QwQ-32B model achieves improvements across all metrics.
@@
We evaluate both LLMs and RLMs on the explanation generation task in TrendFact, focusing on three key aspects: explanation quality, ECS, and HCPI. The results are presented in Table 4. First, we observe that, despite superior performance in the verification task, RLMs generally generate lower-quality explanations compared to LLMs. For example, QwQ-32B, the best performer in verification, produces relatively low-quality explanations. We attribute this to the design of RLMs, which prioritizes concise key information and accurate conclusions over elaborative output, resulting in less detailed and rich explanations. Second, RLMs outperform LLMs in terms of explanation consistency, as measured by ECS, which is expected since ECS captures the alignment between explanations and the model’s internal reasoning process. Furthermore, FactISR improves the explanation quality of RLMs without compromising consistency, enabling them to approach the explanation performance of LLMs. Finally, regarding hotspot perception capability as measured by HCPI, RLMs achieve the best results, and FactISR further enhances their performance in this aspect. Strong hotspot perception requires checking system to consider various event attributes, such as hotspot indicators, importance, and judgment consistency, to effectively identify high-impact samples and reduce the risk of misinformation. Base LLMs often struggle to reason over such complex information, resulting in weaker perception capabilities. In contrast, RLMs leverage their reasoning strengths to capture implicit relationships among attributes to enable perception. Furthermore, FactISR further enhances it by guiding RLMs to reason more effectively.
Methods Acc ECS HCPI
FactISR(QwQ-32B) 74.32 0.8375 0.6336
- w/o DEA 73.30 0.8342 0.6189
- w/o ET 73.16 0.8361 0.6081
- w/o ISR 73.46 0.8328 0.6146
FactISR(Qwen3-32B) 72.46 0.8268 0.6157
- w/o DEA 71.09 0.8168 0.5844
- w/o ET 70.24 0.8230 0.5605
- w/o ISR 71.25 0.8211 0.5935
@@ 5.3 Ablation Study
To evaluate the effectiveness of each component in FactISR, we conduct ablation studies by individually removing one of the three modules from the fully integrated QwQ-32B and Qwen3-32B. We assess performance using both the accuracy of the verification task and the ECS of the explanation generation task. The results are shown in Table 5. It demonstrates that removing any single module leads to performance degradation across both tasks, confirming the effectiveness of all components in FactISR. Furthermore, we conduct further ablation experiments to analyze the efficiency gains introduced by the DEA module. As shown in Table 6, DEA contributes to improved accuracy and substantial gains in reasoning efficiency, achieving average reductions of more than 50% in both reasoning time and length. These results indicate that DEA effectively mitigates the reasoning inefficiencies caused by lengthy evidence.
Methods Acc Time Length
QwQ-32B 72.67 0.8321 2664
+ DEA 73.16 0.4294 ($\downarrow$48.40%) 817 ($\downarrow$69.33%)
Qwen3-32B 70.09 0.7850 2664
+ DEA 70.24 0.3851 ($\downarrow$50.94%) 803 ($\downarrow$69.86%)
@@ 6 Conclusion
In this paper, we introduce TrendFact, the first hotspot perception fact-checking benchmark for evaluation of fact verification, evidence retrieval, and explanation generation tasks. It comprises 7,643 challenging samples and an evidence library containing 66,217 entities through a rigorous data construct process. We also propose two novel metrics, ECS and HCPI, to assess the explanation generation and hotspot perception capability of fact-chekcing systems. In addition, we present FactISR, a method designed to enhance the fact-checking capabilities of RLMs. Experimental results demonstrate that TrendFact poses challenges to numerous advanced LLMs, RLMs, and automated fact-checking methods, while FactISR effectively improves RLMs performance.
@@ Appendix A Details of Data Attributes
The hotspot attributes of TrendFact sample include: views, discussions, posts, and engagements. These indicators are crucial for assessing the hotspot perception capabilities of fact-checking systems. Specifically, views represent the number of times the sample has been viewed on sampled trending platforms; discussions indicate the number of times the sample has been discussed; posts refer to the number of posts triggered by the sample; engagements represent the number of users involved. Additionally, the influence score, which is assessed by an LLM to indicate the potential threat level if the claim were false, ranges from 1 to 5, with 5 being the highest threat. This score is also a key component in evaluating the hotspot perception capabilities of fact-checking systems.
@@
1 You are a trending topics analysis assistant, capable of accurately identifying the category of trending topics, mainly referring to trends on platforms like Weibo and Baidu. I will provide you with data on trending topics, and you need to help me determine their category. Note: I do not want entertainment-related trending topics. This means you do not need to output specific categories; you only need to decide whether a trending topic belongs to the entertainment category, and simply output one word: "Yes" or "No." Next, I will give you several examples for your reference in making judgments and outputs. {Example Trending Topics} Note: To emphasize again, you need to determine if a trending topic belongs to the entertainment category, and output only one word (Yes/No)! Note: If the trending topic is garbled text, also output No!

### v4

@@ 5 FactISR
In this section, we detail our proposed method, FactISR, a reasoning framework designed to enhance the fact-checking capabilities and HPA of RLMs. As shown in Figure 3, it consists of two key components: dynamic evidence augmentation (DEA) and iterative self-reflection (ISR). Unlike traditional RAG, which loads all evidence at once for reasoning, FactISR dynamically calls upon DEA to supplement evidence throughout the reasoning process. And based on the claim’s influence score and the interim reasoning results, ISR is selectively activated by RLM to further refine the process and achieve more accurate outcomes.
Dynamic Evidence Augmentation
Traditional RAG process loads all evidence at once, which can lead to issues such as insufficient, redundant, or irrelevant evidence, hindering accurate judgment. To address these, FactISR’s DEA module dynamically retrieves and adds relevant evidence during the reasoning process based on ongoing analysis of RLMs. This continuous addition of evidence guarantees a more refined evidence and produces better fact-checking performance.
Iterative Self-Reflection
Traditional RAG evaluates all claims equally and lacks the capability to adjust its reasoning budget based on a claim’s popularity. This limitation can lead to redundant inference for low-influence claims and insufficient reasoning for high-influence ones, ultimately detracting from HPA performance. To address these, we introduce the iterative self-reflection module, which enables RLMs to iteratively reflect according to a claim’s popularity.
Specifically, if a claim has a high influence score, the probability of reflection reasoning increases. Moreover, as the reasoning process advances, the probability gradually decreases, effectively preventing excessive reflection. The formal definition of the reflection probability is as follows:
$P_{i}^{n}=(k\cdot{\frac{s_{i}-s_{\text{min}}}{s_{\text{max}}-s_{\text{min}}+\epsilon}})\cdot\gamma^{n-1}$ (4)
Where $P_{i}^{n}$ represents the probability of the $n$-th reflection for the $i$-th sample, while $s_{\text{min}}$ and $s_{\text{max}}$ denote the minimum and maximum scores among all claims, respectively. A small constant $\epsilon$ is included to prevent division by zero, typically set to $1\times 10^{-8}$. The hyperparameter $k\in[0,1]$ is crucial for flexibly controlling the overall reflection intensity across samples, with a default value of 1, and $\gamma\in[0,1]$ is the decay factor. where the superscript $n$ in $\gamma^{n}$ indicates exponentiation.
@@ 6.1 Setup
For evidence retrieval task, we choose R@k, where k=1,2,3,5. For verification task, we choose F1-macro, Precision, Recall, and Accuracy. For explanation generation task, in addition to ECS, we also employ BLEU-4, ROUGE-(1, 2, L), and BERTScore. For assessing the HPA of fact-checking systems, we employ HCPI.
@@
In this work, we choose the following types of methods as baseline, including RLMs, LLMs, and existing fact-checking methods. For RLMs, we select the most advanced QwQ-32B, QwQ-32B-Preview qwe (2024), Qwen3-32B (Think) Yang et al. (2025), and DeepSeek-R1-0528 Guo et al. (2025). For LLMs, we choose GPT-4.1 Hurst et al. (2024), DeepSeek-v3 Liu et al. (2024), and Qwen2.5-72B-Instruct qwe (2024), Qwen3-32B (No Think). For fact-checking methods, we select PROGRAMFC Pan et al. (2023) and CLAIMDECOMP Chen et al. (2022). For retrieve methods, we choose the following advanced methods, including BM-25 (without date), BM-25, OpenAI’s text-embedding-ada-002, and bge-m3 (dense) Chen et al. (2024).
Experimental Settings
The detailed experiment settings are provided in the Appendix J.
@@
BM25 -w/o date 12.08 20.46 25.98 33.70
BM25 12.99 20.76 26.65 34.91
text-emb-ada-002 12.10 21.44 26.88 35.42
bge-m3(dense) 14.10 22.76 29.63 39.97
@@
Methods F1 P R Acc
PROGRAM-FC 40.46 41.66 43.30 43.17
CLAIMDECOMP 42.53 44.18 44.90 45.28
QwQ-32B-Preview 49.52 52.32 53.48 55.29
Qwen2.5-72B-instruct 47.10 53.51 52.51 55.96
Qwen3-32B(No think) 50.06 53.00 53.92 58.84
DeepSeek-V3-0324 52.10 55.19 55.94 60.67
GPT-4.1 52.45 56.23 55.88 61.29
Qwen3-32B(Think) 58.14 58.05 59.82 66.09
QwQ-32B 58.58 58.89 60.09 68.61
DeepSeek-R1-0528 58.89 59.15 60.58 68.44
FactISR(Qwen3-32B) 58.68 58.55 60.84 67.49
FactISR(QwQ-32B) 61.17 61.04 63.37 69.70
@@
QwQ-32B-Preview 0.4820 0.7843 0.1573 0.7525 0.4699 0.2781 0.4048
Qwen2.5-72B-instruct 0.5061 0.7207 0.2632 0.8015 0.5637 0.3712 0.5080
Qwen3-32B(No think) 0.5001 0.7574 0.2491 0.8049 0.5649 0.3658 0.5106
DeepSeek-V3-0324 0.5427 0.7674 0.2360 0.7918 0.5423 0.3427 0.4887
GPT-4.1 0.5488 0.7726 0.2214 0.7735 0.5380 0.3341 0.4736
Qwen3-32B(Think) 0.5575 0.8332 0.2181 0.7836 0.5188 0.3138 0.4603
QwQ-32B 0.5621 0.8531 0.1940 0.7712 0.4896 0.2864 0.4247
DeepSeek-R1-0528 0.5681 0.8399 0.2041 0.7764 0.5016 0.2954 0.4383
FactISR(Qwen3-32B) 0.5926 0.8426 0.2281 0.7912 0.5365 0.3334 0.4833
FactISR(QwQ-32B) 0.6010 0.8610 0.2175 0.7835 0.5156 0.3107 0.4592
@@ 6.2 Main Results
We evaluate the selected retrieval methods based on their ability to retrieve the target gold evidence from the TrendFact evidence library. As shown in Table 3, the best performing method, bge-m3, achieves a relatively low performance with an R@5 less than 40%. This suggests that our unique evidence library construction effectively gathers challenging evidence that distinguishes original gold evidence, thereby increasing the difficulty.
@@
We conduct a comprehensive evaluation of selected methods on the verification task of TrendFact, with the results presented in Table 4. The key findings are as follows: First, traditional fact-checking methods perform the worst, with all verification scores falling below 50%. Second, both LLMs and RLMs perform better than 50%. Notably, RLMs outperform LLMs, as they are better equipped to handle the complex reasoning required by many TrendFact samples. However, even the best-performing DeepSeek-R1-0528 fails to achieve an overall verification score above 60%, highlighting the significant challenge posed by the high-quality and reasoning-intensive nature of TrendFact. Moreover, FactISR contributes to improved verification performance for RLMs. For instance, it enhances the performance of the RLM QwQ-32B, enabling it to achieve an overall verification F1 score of 61, surpassing DeepSeek-R1-0528.
@@
We evaluate both LLMs and RLMs on the explanation generation task in TrendFact, with the results presented in Table 5. Our findings are as follows: First, RLMs generally produce lower-quality explanations compared to LLMs. We attribute this to the design of RLMs, which prioritizes concise key information and accurate conclusions over elaborate output, resulting in less detailed explanations. Second, RLMs outperform LLMs in explanation consistency, as measured by ECS. This is expected, as ECS reflects the alignment between explanations and the model’s internal reasoning process. Furthermore, FactISR enhances both the explanation quality and consistency of RLMs. For example, FactISR improves the performance of QwQ-32B by 1-3 percentage points across all explanation metrics, enabling it to approach the performance of LLMs in explanation quality.
HPA Assessment Results
Since the verification results affect the calculation of HCPI, the RLM naturally exhibits a higher HCPI compared to the LLM. However, because the RLM lacks the capability to adjust the reasoning budget according to the claim’s influence, it may result in redundant inference for low-influence claims and insufficient reasoning for high-influence claims, adversely affecting its HPA performance. In contrast, FactISR effectively addresses these issues by leveraging the influence score-based ISR module combined with DEA, and improves the RLM’s HCPI score by up to nearly 4%.
Methods F1 ECS HCPI
FactISR(QwQ-32B) 61.17 0.8610 0.6010
- w/o DEA 59.79 0.8501 0.5833
- w/o ISR 58.88 0.8388 0.5675
FactISR(Qwen3-32B) 58.68 0.8426 0.5926
- w/o DEA 58.46 0.8372 0.5785
- w/o ISR 58.30 0.8390 0.5650
@@ 6.3 Ablation Study
Component Effectiveness
We conduct experiments by individually removing one of the two modules from the fully integrated QwQ-32B and Qwen3-32B. We evaluate comprehensive performance using accuracy, ECS, and HCPI. The results are shown in Table 6. It demonstrates that removing any single module leads to performance degradation across all tasks, confirming the effectiveness of all components in FactISR.
DEA Efficiency
We conduct experiments to analyze the efficiency gains introduced by the DEA module. As shown in Table 7, it contributes to maintaining accuracy while achieving substantial gains in reasoning efficiency, with maximum reductions of 30% and 50% in reasoning time and length, respectively. These results indicate that DEA effectively mitigates the reasoning inefficiencies caused by lengthy evidence.
Evaluation with Limited Resources
We conduct experiments to evaluate the influence of FactISR on the HPA performance and fact-checking efficiency of RLMs under resource constraints, as shown in Figure 4. We adjust the number of input evidence for RAG and the reflection intensity hyperparameter $k$ for FactISR. It can be seen that FactISR achieves higher HPA with the same amount of evidence and significantly reduces the average inference time required to reach the same HPA, thereby greatly improving verification efficiency.
Methods F1 Time Length
QwQ-32B 58.89 0.5960 2664
+ DEA 58.88 0.4594 ($\downarrow$22.92%) 1442 ($\downarrow$47.31%)
Qwen3-32B 58.14 0.5515 2664
+ DEA 58.30 0.4380 ($\downarrow$30.17%) 1316 ($\downarrow$50.62%)
@@
Figure 4: Performance comparison between RAG and FactISR under resource constraints.
@@ 7 Conclusion
In this paper, we introduce TrendFact, the first hotspot perception fact-checking benchmark for the evaluation of fact verification, evidence retrieval, and explanation generation tasks. It comprises 7,643 challenging samples and an evidence library containing 366,634 entities through a rigorous data construct process. We also propose two novel metrics, ECS and HCPI, to assess the explanation generation and hotspot perception capability of fact-chekcing systems. In addition, we present FactISR, a method designed to enhance the fact-checking capabilities of RLMs. Experimental results demonstrate that TrendFact poses challenges to numerous advanced LLMs, RLMs, and automated fact-checking methods, while FactISR effectively improves performance.
@@ 8 Limitations
9 Ethics and Compliance
The claims in TrendFact are derived from trending platforms and existing datasets, with evidence sourced from publicly content on the Internet, and its construction does not violate relevant platform rules or legal regulations Calzada (2022); Chen and Sun (2021), making it suitable for academic research. Specifically, the claims collected from trending platforms are taken exclusively from public pages that do not require login. The collection process strictly follows the platform’s open-access agreements (e.g., Weibo Open Platform Agreement: https://open.weibo.com), and only public data is retrieved, without involving any non-public interfaces, private user content, or data requiring authorization. During it, we only collected information such as trending titles, dates, and urls, which fall within the lowest-risk category, and do not include sensitive data such as user nicknames or user IDs. Date information is retained only to the level of “year–month–day” rather than exact timestamps; additionally, we applied rate limiting ($\leq 1$ request per second) to avoid imposing load on the servers.
@@ Appendix A Details of Data Attributes
The hotspot indicators of the TrendFact sample include: views, discussions, posts, and engagements. These indicators are crucial for assessing the hotspot perception capabilities of fact-checking systems. Specifically, views represent the number of times the sample has been viewed on sampled trending platforms; discussions indicate the number of times the sample has been discussed; posts refer to the number of posts triggered by the sample; engagements represent the number of users involved. Additionally, the influence score, which is assessed by an LLM to indicate the potential threat level if the claim were false, ranges from 1 to 5, with 5 being the highest threat. This score is also a key component in evaluating the hotspot perception capabilities of fact-checking systems. Figure 5 shows the data distribution of TrendFact samples, including labels, gold evidence count, and domains.
Figure 5: Overview of the data distribution, including labels, gold evidence count, and domains.
@@
1
You are a trending topics analysis assistant, capable of accurately identifying the category of trending topics, mainly referring to trends on platforms like Weibo and Baidu. I will provide you with data on trending topics, and you need to help me determine their category. Note: I do not want entertainment-related trending topics. This means you do not need to output specific categories; you only need to decide whether a trending topic belongs to the entertainment category, and simply output one word: "Yes" or "No." Next, I will give you several examples for your reference in making judgments and outputs. {Example Trending Topics} Note: To emphasize again, you need to determine if a trending topic belongs to the entertainment category, and output only one word (Yes/No)! Note: If the trending topic is garbled text, also output No!